OpenStudy (anonymous):

Can you solve this for me? sin((pi/6)+x)=cos((pi/4)-x)

myininaya (myininaya):

\[\sin(x)=\cos(\frac{\pi}{2}-x) ; \cos(x)=\sin(\frac{\pi}{2}-x)\] \[\sin(\frac{\pi}{6}+x)=\sin(\frac{\pi}{2}-(\frac{\pi}{4}-x))\] \[\cos(\frac{\pi}{4}-x)=\cos(\frac{\pi}{2}-(\frac{\pi}{6}+x))\]

myininaya (myininaya):

i used those two top equations to rewrite what you have

myininaya (myininaya):

no I suppose we should put something in there like 2npi

myininaya (myininaya):

\[\sin(\frac{\pi}{6}+x)=\sin(\frac{\pi}{2}-(\frac{\pi}{4}-x)+2n \pi)\] \[\cos(\frac{\pi}{4}-x)=\cos(\frac{\pi}{2}-(\frac{\pi}{6}+x)+2n \pi)\]

myininaya (myininaya):

where n is an integer

myininaya (myininaya):

now to solve both equations set the insides equal and solve for x

OpenStudy (anonymous):

Thanks a lot:)

myininaya (myininaya):

:)

OpenStudy (mertsj):

\[\sin(\frac{\pi}{6}+x)=\cos(\frac{\pi}{4}-x)\]

OpenStudy (mertsj):

\[\sin\frac{\pi}{6}cosx+\cos\frac{\pi}{6}sinx=\cos\frac{\pi}{4}cosx+\sin\frac{\pi}{4}sinx\]

OpenStudy (mertsj):

\[\frac{1}{2}cosx+\frac{\sqrt{3}}{2}sinx=\frac{\sqrt{2}}{2}cosx+\frac{\sqrt{2}}{2}sinx\]

OpenStudy (mertsj):

\[(\frac{\sqrt{3}-\sqrt{2}}{2})sinx=(\frac{\sqrt{2}-1}{2})cosx\]

OpenStudy (mertsj):

\[sinx=(\frac{\sqrt{2}-1}{\sqrt{3}-\sqrt{2}})cosx\]

OpenStudy (mertsj):

\[tanx=\frac{\sqrt{2}-1}{\sqrt{3}-\sqrt{2}}\]

OpenStudy (mertsj):

\[x=\frac{7\pi}{24}\]